(1) Your patient Mamine is a 15-year-old with end-stage renal disease. She has Vancomycin 440 mg IV piggyback once. You have vancomycin 1 g recommended with 20 mL of sterile water. How many milliliters should you add to the 100 mL minibag to administer the 440 mg dose

(2) Betony is a 2 ½ year old who is newly diagnosed with HIV infection. She is to be started on Retrovir 120 mg by mouth every 12 hours. She weighs 13 kg and is 90 cm long.

  1. Using the formula for BSA, calculate her body surface area

BSA (mu2) – Ht (cm) x Wt (Kg) / 3600

  1. Retrovir is prescribed on an individual basis ranging from 180 to 240 mg/m2 per dose every 12 hours. It comes in a concentration of 10 mg/mL.

Is her dose of 120 mg within the acceptable range?

How many mL will you administer?
Answer #1

1) Recommended dose = 1 gm/20ml (1m = 1000mg) Ordered dose = 440mg. If logoing in 2001 Then 440 mg in xml lovo xz = 20x440z ] ay ordered dose-120mg weight = 13kg Height =90cm : Formula = BSA (m²) = H1(cm) xwt (kg) + 3600 = 90 x 13 = 3600 - 2170 368 Available dose = long 1ml ordered dose = 120mg If 10 mg in Imlu Then 120 mg in xml a 10xx = 1 X 120 n. 2 = 120 - 12 m 10 ..